# Invariant subspace perturbations related to defective eigenvalues of $\Delta$-Hermitian and Hamiltonian matrices

### v1

@@ 3.2 Real eigenvalue case
for $k=1,2,\ldots,m$. ($\widehat{W}_{m+1}$ and $E_{m+1}$ are void.) Note $\widehat{W}_{k}$ is Hermitian for any $k$.
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
@@
Under the conditions of Theorem 6, suppose $F$ is a full column rank matrix of the form (84) that satisfies (83). and $\Omega$ contains part of the eigenvalues of $\Theta_{\rho}(S_{\rho})$ that are either real or in complex conjugate pairs, and $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset$. Then for $t$ sufficiently small, there exist matrices $\widetilde{F}(t)$ and $\widetilde{\Omega}(t)$ satisfying (85) and (86), such that the matrix
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu(q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@ 4.1 Nonimaginary eigenvalue case
Since $\imath J_{n}$ and $\Delta_{T}$ are congruent to $\mathop{\rm diag}\nolimits(I_{n},-I_{n})$ and $\mathop{\rm diag}\nolimits(I_{p},-I_{p})$, respectively, $\Delta_{C}$ in (92) is congruent to $\mathop{\rm diag}\nolimits(I_{n-p},-I_{n-p})$. Therefore, there is $\widehat{P}_{C}$ such that
@@
Let $H$ and $T$ be Hamiltonian matrices and $H$ satisfy (95) with $\lambda$ being nonimaginary, $\Phi=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right]$, $\Phi_{T}=\left[\begin{array}[]{cc}\Xi&\Xi_{c}\end{array}\right]$. Suppose $t$ is sufficiently small. Let $\Psi_{1},\ldots,\Psi_{m}$ be defined in (98). For a fixed $\rho\in\{1,\ldots,m\}$ assume $\Psi_{\rho+1}$ is invertible if $\rho<m$. and let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix $\widehat{S}_{\rho}$, where $\widehat{S}_{\rho}$ is defined in (102).
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
@@
$\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad(\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\widetilde{F}^{(c)}(t)=\widetilde{F}^{(c)}(t)\widetilde{\Omega}^{\ast}(t),\qquad\widetilde{F}^{\ast}\widetilde{F}^{(c)}=I_{\ell}.$
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),k>\rho\end{array}\right$.
@@
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.

### v2

@@ 3.2 Real eigenvalue case
for $k=1,2,\ldots,m-1$, and $\widehat{W}_{m}=U_{m1}^{\ast}\Delta DU_{m1}$. Note $\widehat{W}_{k}$ is Hermitian for any $k$.
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
Under the conditions of Theorem 8, suppose $F$ is a full column rank matrix of the form (84) that satisfies (83). Also, suppose $\Lambda(\Omega)\subseteq\Lambda(\Theta_{\rho}(S_{\rho}))$ with the eigenvalues in $\Lambda(\Omega)$ being either real or in complex conjugate pairs and $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset$. Then for $t$ sufficiently small, there exist matrices $\widetilde{F}(t)$ and $\widetilde{\Omega}(t)$ satisfying (85) and (86), such that the matrix
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu(q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@ 4.1 Nonimaginary eigenvalue case
Since $\imath J_{n}$ and $\Delta_{T}$ are congruent to $\mathop{\rm diag}\nolimits(I_{n},-I_{n})$ and $\mathop{\rm diag}\nolimits(I_{p},-I_{p})$, respectively, $\Delta_{C}$ in (92) is congruent to $\mathop{\rm diag}\nolimits(I_{n-p},-I_{n-p})$. Therefore, there is a $\widehat{P}_{C}$ such that
@@
Let $H$ and $T$ be Hamiltonian matrices and $H$ satisfy (95) with $\lambda$ being nonimaginary, $\Phi=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right]$, $\Phi_{T}=\left[\begin{array}[]{cc}\Xi&\Xi_{c}\end{array}\right]$. Suppose $t$ is sufficiently small. Let $\Psi_{1},\ldots,\Psi_{m}$ be defined in (98). For a fixed $\rho\in\{1,\ldots,m\}$ assume $\Psi_{\rho+1}$ is invertible if $\rho<m$, and let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix $\widehat{S}_{\rho}$, where $\widehat{S}_{\rho}$ is defined in (102).
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
$\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad(\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\widetilde{F}^{(c)}(t)=\widetilde{F}^{(c)}(t)\widetilde{\Omega}^{\ast}(t),\qquad\widetilde{F}^{\ast}\widetilde{F}^{(c)}=I_{\ell},$
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.